# Efficient Segmentation of Search Markets

### v2

@@ 3.  First-best market segmentation
To characterize the solution to (FB-$\sigma$), let the inverse of the mapping $t\mapsto 1/m^{\prime}(t)$ be denoted by $f$ so that $f(y)=t$ if $1/m^{\prime}(t)=y$. Since $m^{\prime}(t)$ is continuous and strictly decreasing in $t$, $f(y)$ is continuous and strictly increasing in $y$. Furthermore, under Assumption 1, $\lim_{t\to 0}m^{\prime}(t)=\beta$ and $\lim_{t\to\infty}m^{\prime}(t)=0$, [^9] Since $m$ is strictly concave and $m(0)=0$ ( Assumption 1), $m^{\prime}(t)<m(t)/t$ for all $t>0$. Taking limits on both sides as $t\to 0$, we have $m^{\prime}(0)\leq\beta$. Similarly, given $t>0$, we have $m^{\prime}(t)>\frac{m(t^{\prime})-m(t)}{t^{\prime}-t}$ for all $t^{\prime}>t$. Taking double limits on both side, first as $t\to 0$ and then as $t^{\prime}\to 0$, yields that $m^{\prime}(0)\geq\beta$. which implies that $f:[1/\beta,\infty)\to\mathbb{R}_{+}$ with $f(1/\beta)=0$ and $\lim_{y\to\infty}f(y)=\infty$.
@@
$k\int_{\Delta(\Theta)}f\left(\frac{\mathbb{E}_{\nu}[\theta]}{\eta}\right)\cdot\mathbbm{1}_{\left[\mathbb{E}_{\nu}[\theta]>\frac{\eta}{\beta}\right]}\sigma_{\mu}(d\nu)=1.$
@@
$\tau^{FB}_{\sigma}(\nu)=\left\{\begin{array}[]{ccc}0&\mbox{if}&\mathbb{E}_{\nu}[\theta]\leq\eta_{\sigma}/\beta\\[4.0pt] \displaystyle f\left(\frac{\mathbb{E}_{\nu}[\theta]}{\eta_{\sigma}}\right)&\mbox{if}&\mathbb{E}_{\nu}[\theta]>\eta_{\sigma}/\beta\end{array}\right.$
@@
Perfect segmentation $\sigma^{PS}$ is a first-best market segmentation.
@@ 4.  Search Equilibrium
A search equilibrium of a market segmentation $\sigma$ is given by a submarket tightness function $\tau:\Delta(\Theta)\to\mathbb{R}_{+}$ and a reservation value $u\in\mathbb{R}$ such that the pair $(\tau,u)$ jointly satisfies
@@
for $\sigma_{\mu}$-almost all $\nu$ with equality if $\tau(\nu)>0$, and
$u=k\int_{\Delta(\Theta)}m(\tau(\nu))\mathbb{E}_{\nu}[\lambda(\theta)\theta]\sigma_{\mu}(d\nu).$ (4)
In essence, buyers form expectations about submarket tightness $\tau$ and their reservation value $u$—the average payoff they expect to earn from participating in the search market. In equilibrium, these expectations must be consistent with the actual outcomes of buyers’ optimal search behavior. Specifically, buyers adapt a cutoff strategy: submarkets in which the buyers’ expected payoff falls below the reservation value are inactive, while those offering payoffs at least as large as the reservation value are active. Furthermore, among active submarkets, buyers must be indifferent between any two; otherwise, they would all gravitate toward the submarket offering the higher expected payoff. This rational expectations condition regarding submarket tightness is formally expressed in (1).
Additionally, buyers’ reservation value must also be consistent their average payoff generated from trade across all active submarket. This rational expectations condition regarding the payoff is captured by (4).
Notice that Definition 1 does not explicitly require $\tau$ to satisfy the feasibility constraint given by (3). However, if a pair $(\tau,u)$ is a search equilibrium of market segmentation $\sigma$, then $\tau$ necessarily satisfies (3). To see why, notice that (1) implies that $m(\tau(\nu))\mathbb{E}_{\nu}[\lambda(\theta)\theta]=u\tau(\nu)$ for $\sigma_{\mu}$-almost all $\nu$. Integrating over all submarkets then yields
$\int_{\Delta(\Theta)}m(\tau(\nu))\mathbb{E}_{\nu}[\lambda(\theta)\theta]\sigma_{\mu}(d\nu)=u\int_{\Delta(\Theta)}\tau(\nu)\sigma_{\mu}(d\nu).$
Combining the above equality with (4) then immediately implies that (3) is satisfied.
@@
$\tau^{*}_{\sigma}(\nu)=\left\{\begin{array}[]{ccc}0&\mbox{if}&\mathbb{E}_{\nu}[\lambda(\theta)\theta]\leq u^{*}_{\sigma}/\beta\\[4.0pt] \displaystyle g\left(\frac{\mathbb{E}_{\nu}[\lambda(\theta)\theta]}{u^{*}_{\sigma}}\right)&\mbox{if}&\mathbb{E}_{\nu}[\lambda(\theta)\theta]>u^{*}_{\sigma}/\beta\end{array}\right..$
@@
To build intuition, consider a fixed reservation value $u$. When buyers anticipate a payoff of $u$, they optimally avoid entering any submarket $\nu$ satisfying $\beta\cdot\mathbb{E}_{\nu}[\lambda(\theta)\theta]\leq u$, even if these submarkets offer the highest probability of meeting a seller. This is because the expected payoff from trade in such submarkets is too low to justify participation. Among the remaining submarkets, those offering a higher payoff conditional on trade attract more buyers, leading to greater congestion. As a result, congestion in each active submarket precisely offsets its potential gains from trade. Consequently, the optimal submarket tightness for a given reservation value $u$ is determined by the mapping $\nu\mapsto g\left(\frac{\mathbb{E}_{\nu}[\lambda(\theta)\theta]}{u}\right)\mathbbm{1}_{[\beta\cdot\mathbb{E}_{\nu}[\lambda(\theta)\theta]>u]}$. Moreover, a buyer’s average payoff across all active submarkets is given by
$\int_{\Delta(\Theta)}m\left(g\left(\frac{\mathbb{E}_{\nu}[\lambda(\theta)\theta]}{u}\right)\right)\mathbbm{1}_{[\beta\cdot\mathbb{E}_{\nu}[\lambda(\theta)\theta]>u]}\mathbb{E}_{\nu}[\lambda(\theta)\theta]\sigma_{\mu}(d\nu),$
@@ 5.1.  Hosios condition
Despite the potential inefficiencies in the equilibrium allocation of buyers, this section identifies necessary and sufficient conditions under which the search equilibrium of a perfectly segmented market constitutes a first-best outcome, as established in Proposition 1 and Proposition 2. To state the result, let $\varepsilon:\mathbb{R}_{+}\to\mathbb{R}_{+}$ denote the elasticity of the meeting function with $\varepsilon(t)\coloneqq t\cdot m^{\prime}(t)/m(t)$ for $t\geq 0$.
@@ 5.2.  Constrained-efficient market segmentation
This optimization problem is inherently complex: it involves optimizing over an infinite-dimensional choice variable that enters the objective function in a highly non-linear way. Specifically, the choice of segmentation influences buyers’ beliefs about seller types within each submarket, which in turn affects how buyers allocate themselves across submarkets. [^10] Although (5.2) resembles an information design problem, it differs in a crucial way: whereas a receiver’s optimal action in standard information design depends solely on posterior beliefs, buyers’ search strategies here depend both on their posterior beliefs as well as the entire distribution of posterior beliefs. In Section A.2, I derive a sufficient and necessary condition that decouples these two interdependent effects. However, even this requires solving a constrained concavification problem (Kamenica and Gentzkow, 2011) with a continuum of states. As such, a general characterization of the constrained-efficient market segmentation appears intractable in many cases.
@@
To that end, let $F:\Theta\to[0,1]$ denote the cumulative distribution function (CDF) associated with the prior $\mu\in\Delta(\Theta)$. Importantly, $F$ is also the posterior-mean distribution induced by the perfect market segmentation $\sigma^{PS}$, with
$F(x)\coloneqq\sigma^{PS}_{\mu}\big{(}\{\nu\in\Delta(\Theta):\mathbb{E}_{\nu}[\theta]\leq x\}\big{)}$
for all $x\in\Theta$. More generally, a CDF $H:\Theta\to[0,1]$ represents a posterior-mean distribution induced by some market segmentation if and only if $H$ is a mean-preserving contraction of $F$ (Blackwell, 1953; Blackwell and Girshick, 1954). Let $\operatorname*{MPC}(F)$ denote the set of all such mean-preserving contractions.
@@
A market segmentation $H\in\operatorname*{MPC}(F)$ is constrained-efficient if and only if
@@
Furthermore,
  1. ()()( )
  2. ()()( )
    $H(x)=\left\{\begin{array}[]{ccc}0&\mbox{if}&x<\underline{x}\\ F(\theta^{c})&\mbox{if}&x\in[\underline{x},\bar{x})\\ 1&\mbox{if}&x\geq\bar{x}\\ \end{array}\right.$
    is a constrained-efficient segmentation. In this case, $\tau_{H}(\underline{x})=0<\tau_{H}(\bar{x})$.
@@ A.1.  Proofs
$\psi(\nu,\eta)\coloneqq\left\{\begin{array}[]{ccc}0&\mbox{if}&\mathbb{E}_{\nu}[\theta]\leq\eta/\beta\\ f\left(\frac{\mathbb{E}_{\nu}[\theta]}{\eta}\right)&\mbox{if}&\mathbb{E}_{\nu}[\theta]>\eta/\beta\end{array}\right..$
@@ 4. $(iv)$
Fix any market segmentation $\sigma$, and let $\tau^{FB}_{\sigma}$ be the solution to (FB-$\sigma$). The proof proceeds by showing that, given the pair $(\sigma,\tau_{\sigma}^{FB})$, there exists another pair $(\sigma^{PS},\tilde{\tau})$ that achieves a (weakly) higher surplus. To that end, define a new submarket tightness function $\tilde{\tau}:\Delta(\Theta)\to\mathbb{R}_{+}$ given by
@@
where the first equality follows from the definition of $\sigma^{PS}$ and the second equality follows by construction of $\tilde{\tau}$. Therefore, $\tilde{\tau}\in\mathcal{L}^{1}(\Delta(\Theta),\sigma^{PS}_{\mu})$ and the pair $(\sigma^{PS},\tilde{\tau})$ satisfies (3).
Evaluating the expected surplus under $(\sigma,\tau_{\sigma})$, we have
$k\int_{\Delta(\Theta)}m(\tau^{FB}_{\sigma}(\nu))\mathbb{E}_{\nu}[\theta]\sigma_{\mu}(d\nu)=$ $k\int_{\Theta}\int_{\Delta(\Theta)}m(\tau^{FB}_{\sigma}(\nu))\theta\sigma(d\nu,\theta)\mu(d\theta)$
$\leq$ $k\int_{\Theta}m\left(\int_{\Delta(\Theta)}\tau^{FB}_{\sigma}(\nu)\sigma(d\nu,\theta)\right)\theta\mu(d\theta)$
$=$ $k\int_{\Theta}m(\tilde{\tau}(\delta_{\theta}))\theta\mu(d\theta)$
$=$ $k\int_{\Delta(\Theta)}m(\tilde{\tau}(\nu))\mathbb{E}_{\nu}[\theta]\sigma^{PS}_{\mu}(d\nu),$
where the first equality follows from changing the order of integration and the fact that $\sigma$ satisfies (2.1), the inequality follows from the concavity of $m$, the second equality follows by definition of $\tilde{\tau}$, and the last follows by the definition of $\sigma^{PS}$.
Thus, for any arbitrary market segmentation $\sigma$, the pair $(\sigma,\tau_{\sigma}^{FB})$ yields a (weakly) lower surplus than the pair $(\sigma^{PS},\tilde{\tau})$. Of course, $(\sigma^{PS},\tilde{\tau})$ yields a (weakly) lower surplus than the pair $(\sigma^{PS},\tau^{FB}_{\sigma^{PS}})$, because the latter solves (FB-$\sigma$) when $\sigma=\sigma^{PS}$. As a result, the pair $(\sigma^{PS},\tau^{FB}_{\sigma^{PS}})$ is a solution to (FB). ∎
@@
where the equality follows from (1), and the inequality follows from the fact $\beta$ is the highest probability with which a buyer meets a seller.
@@
where the inequality follows from (1), and the equality follows by definition that $m(0)/0=\beta$. Thus, from the contrapositive, $\beta\mathbb{E}_{\nu}[\lambda(\theta)\theta]>u^{*}_{\sigma}$ implies $\tau^{*}_{\sigma}(\nu)>0$. ∎
Let us now turn to proving the existence of an essentially-unique equilibrium. By Lemma 1, any search equilibrium $(\tau^{*}_{\sigma},u^{*}_{\sigma})$ satisfies $\tau^{*}_{\sigma}(\nu)=0$ for $\sigma_{\mu}$-almost all $\nu$ such that $\beta\mathbb{E}_{\nu}[\lambda(\theta)\theta]\leq u^{*}_{\sigma}$, and $\tau^{*}_{\sigma}(\nu)>0$ for $\sigma_{\mu}$-almost all $\nu$ such that $\beta\mathbb{E}_{\nu}[\lambda(\theta)\theta]>u^{*}_{\sigma}$. In the latter case, (1) implies that
@@
which from the definition of $g$ implies that $\tau^{*}_{\sigma}(\nu)=g\left(\mathbb{E}_{\nu}[\lambda(\theta)\theta]/u^{*}_{\sigma}\right)$. In other words, there exists a mapping ϕ:Δ(Θ)×(0,)+:italic-$\phi:\Delta(\Theta)\times(0,\infty)\to\mathbb{R}_{+}$ given by
$\phi(\nu,u)\coloneqq\left\{\begin{array}[]{ccc}0&\mbox{if}&\mathbb{E}_{\nu}[\lambda(\theta)\theta]\leq u/\beta\\[4.0pt] \displaystyle g\left(\frac{\mathbb{E}_{\nu}[\lambda(\theta)\theta]}{u}\right)&\mbox{if}&\mathbb{E}_{\nu}[\lambda(\theta)\theta]>u/\beta\end{array}\right.$
@@
Next, observe that a pair $(\tau^{*}_{\sigma},u^{*}_{\sigma})$ jointly satisfies (1) and (4) if and only if $u^{*}_{\sigma}$ is a fixed point of the function $\Phi:(0,\infty)\to\mathbb{R}_{+}$ given by
@@
Thus far, we have shown that $\Phi(u)$ is a continuous and (weakly) decreasing function with $\lim_{u\to\beta\sup_{\theta\in\Theta}\lambda(\theta)\theta}\Phi(u)=0<\lim_{u\to 0}\Phi(u)$. Therefore, it has a unique fixed point in the interval $(0,\beta\sup_{\theta\in\Theta}\lambda(\theta)\theta)$. We have thus established the existence of an essentially-unique pair $(\tau^{*}_{\sigma},u^{*}_{\sigma})$ with $u^{*}_{\sigma}=\Phi(u^{*}_{\sigma})$ and $\tau^{*}_{\sigma}(\cdot)=\phi(\cdot,u^{*}_{\sigma})$, such that $(\tau^{*}_{\sigma},u^{*}_{\sigma})$ jointly satisfies (1) and (4). In other words, there exists an essentially-unique search equilibrium $(\tau^{*}_{\sigma},u^{*}_{\sigma})$ for each market segmentation $\sigma$. ∎
@@
Thus, $\lambda(\theta)\theta\leq\lambda(\frac{\eta_{\sigma^{PS}}}{\beta})\frac{\eta_{\sigma^{PS}}}{\beta}$ for all $\theta\leq\eta_{\sigma^{PS}}/\beta$.
@@
(If direction): Suppose the surplus-splitting function, $\lambda$, satisfies the properties given in the proposition. We first show that $(\tau^{FB}_{\sigma^{PS}},\lambda(\frac{\eta_{\sigma^{PS}}}{\beta})\eta_{\sigma^{PS}})$ is a search equilibrium of $\sigma^{PS}$.
@@
where the first equality follows from the characterization of $\tau^{FB}_{\sigma^{PS}}(\delta_{\theta})$ for $\theta>\eta_{\sigma^{PS}}/\beta$ (Proposition 1), the second equality follows by the assumption on $\lambda$ given in the proposition, and the last equality follows from the fact that $f$ is the inverse of the mapping $t\mapsto 1/m^{\prime}(t)$. Therefore, the pair $(\tau^{FB}_{\sigma^{PS}},\lambda(\eta_{\sigma^{PS}})\eta_{\sigma^{PS}})$ satisfies (1) of Definition 1.

### v3

@@ 3.  First-best market segmentation
To characterize the solution to (FB-$\sigma$), let the inverse of the mapping $t\mapsto 1/m^{\prime}(t)$ be denoted by $f$ so that $f(y)=t$ if $1/m^{\prime}(t)=y$. Since $m^{\prime}(t)$ is continuous and strictly decreasing in $t$, $f(y)$ is continuous and strictly increasing in $y$. Furthermore, under Assumption 1, $\lim_{t\to 0}m^{\prime}(t)=\beta$ and $\lim_{t\to\infty}m^{\prime}(t)=0$, [^8] Since $m$ is strictly concave and $m(0)=0$ ( Assumption 1), $m^{\prime}(t)<m(t)/t$ for all $t>0$. Taking limits on both sides as $t\to 0$, we have $m^{\prime}(0)\leq\beta$. Similarly, given $t>0$, we have $m^{\prime}(t)>(m(t^{\prime})-m(t))/(t^{\prime}-t)$ for all $t^{\prime}>t$. Taking double limits on both side, first as $t\to 0$ and then as $t^{\prime}\to 0$, yields that $m^{\prime}(0)\geq\beta$. which implies that $f:[1/\beta,\infty)\to\mathbb{R}_{+}$ with $f(1/\beta)=0$ and $\lim_{y\to\infty}f(y)=\infty$.
@@
$k\int_{\Delta(\Theta)}f\left(\frac{\mathbb{E}_{\nu}[\theta]}{\eta}\right)\cdot\mathbbm{1}_{\left[\beta\,\mathbb{E}_{\nu}[\theta]>\eta\right]}\sigma_{\mu}(d\nu)=1.$
@@
$\tau^{FB}_{\sigma}(\nu)=\left\{\begin{array}[]{ccc}0&\mbox{if}&\beta\,\mathbb{E}_{\nu}[\theta]\leq\eta_{\sigma}\\[10.0pt] \displaystyle f\left(\frac{\mathbb{E}_{\nu}[\theta]}{\eta_{\sigma}}\right)&\mbox{if}&\beta\,\mathbb{E}_{\nu}[\theta]>\eta_{\sigma}\end{array}\right$.
@@
To characterize the first-best market segmentation, I begin by defining the notion of a perfect market segmentation, which allocates each seller type to a distinct submarket. Formally, for each $\theta\in\Theta$, let $\delta_{\theta}\in\Delta(\Theta)$ denote the Dirac measure at $\theta$, so that for any $A\in\mathcal{B}(\Theta)$, $\delta_{\theta}(A)=\mathbbm{1}_{[\theta\in A]}$. The perfect market segmentation is then defined by the stochastic kernel $\sigma^{PS}$ such that, for all $B\in\mathcal{B}(\Delta(\Theta))$ and $\theta\in\Theta$, $\sigma^{PS}(B,\theta)=\mathbbm{1}_{[\delta_{\theta}\in B]}$.
@@
The perfect segmentation $\sigma^{PS}$ is a first-best market segmentation.
@@ 4.  Search Equilibrium
A search equilibrium of a market segmentation $\sigma$ is given by a submarket tightness function $\tau:\Delta(\Theta)\to\mathbb{R}_{+}$ and a reservation value $u\in\mathbb{R}$ such that
  1. $(a)$
    $\tau$ satisfies the feasibility constraint (3), and
  2. $(b)$
    $(\tau,u)$ jointly satisfy
@@ 2. $(b)$
    for $\sigma_{\mu}$-almost all $\nu$ with equality if $\tau(\nu)>0$.
In essence, buyers form expectations about submarket tightness $\tau$ and their reservation value $u$, which represents the average payoff they expect to earn from participating in the search market. In equilibrium, these expectations must be consistent with a feasible allocation of buyers across submarkets, as captured by (3), as well as with an optimal buyers’ search strategy, as captured by (b). Intuitively, the latter condition states that submarkets that yield a payoff lower than $u$ remain inactive, while those offering payoffs at least as large as $u$ are active. Furthermore, among active submarkets, buyers must be indifferent between any two; otherwise, they would all gravitate toward the submarket offering the higher expected payoff.
Finally, a search equilibrium $(\tau,u)$ must satisfy a rational expectations condition: the anticipated reservation value $u$ should be consistent with the buyers’ actual ex-ante payoff. To see this, observe that a pair $(\tau,u)$ that satisfies Definition 1 implies that $m(\tau(\nu))\mathbb{E}_{\nu}[\lambda(\theta)\theta]=u\,\tau(\nu)$ for $\sigma_{\mu}$-almost all $\nu$ from (b). Integrating over all submarkets then yields
$\int_{\Delta(\Theta)}m(\tau(\nu))\mathbb{E}_{\nu}[\lambda(\theta)\theta]\sigma_{\mu}(d\nu)$ $=u\int_{\Delta(\Theta)}\tau(\nu)\sigma_{\mu}(d\nu)$
$=\frac{u}{k},$
where the last equality follows from (3). On the other hand, buyers’ ex-ante payoff from an allocation policy $Q\in\Delta(\Delta(\Theta))$ with $dQ/d\sigma_{\mu}=k\cdot\tau$ is given by [^9] Recall that given segmentation $\sigma$, submarket $\nu\in\operatorname*{supp}(\sigma_{\mu})$ has a measure $k\cdot\sigma_{\mu}(d\nu)$ sellers. Hence, if the submarket has tightness $\tau(d\nu)$, then it must have a measure $Q(d\nu)$ of buyers, where $Q(d\nu)=k\cdot\tau(d\nu)\cdot\sigma_{\mu}(d\nu)$.
$\int_{\Delta(\Theta)}U(\nu;\tau)Q(d\nu)$ $=k\int_{\Delta(\Theta)}m(\tau(\nu))\mathbb{E}_{\nu}[\lambda(\theta)\theta]\sigma_{\mu}(d\nu).$
Putting these two expressions together then yields the desired rational expectations condition:
$u=k\int_{\Delta(\Theta)}m(\tau(\nu))\mathbb{E}_{\nu}[\lambda(\theta)\theta]\sigma_{\mu}(d\nu),$ (4)
that is, the buyers’ reservation value is equal to the share of the ex-ante surplus they capture in equilibrium.
@@
$\tau^{*}_{\sigma}(\nu)=\left\{\begin{array}[]{ccc}0&\mbox{if}&\beta\,\mathbb{E}_{\nu}[\lambda(\theta)\theta]\leq u^{*}_{\sigma}\\[4.0pt] \displaystyle g\left(\frac{\mathbb{E}_{\nu}[\lambda(\theta)\theta]}{u^{*}_{\sigma}}\right)&\mbox{if}&\beta\,\mathbb{E}_{\nu}[\lambda(\theta)\theta]>u^{*}_{\sigma}\end{array}\right.$.
@@
To build intuition, consider a fixed reservation value $u$. When buyers anticipate a payoff of $u$, they optimally avoid entering any submarket $\nu$ satisfying $\beta\cdot\mathbb{E}_{\nu}[\lambda(\theta)\theta]\leq u$, even if these submarkets offer the highest probability of meeting a seller. This is because the expected payoff from trade in such submarkets is too low to justify participation. Among the remaining submarkets, those offering a higher payoff conditional on trade attract more buyers, leading to greater congestion. As a result, congestion in each active submarket precisely offsets its potential gains from trade. Consequently, the optimal submarket tightness for a given reservation value $u$ is determined by the mapping
$\nu\mapsto g\left(\frac{\mathbb{E}_{\nu}[\lambda(\theta)\theta]}{u}\right)\mathbbm{1}_{[\beta\,\mathbb{E}_{\nu}[\lambda(\theta)\theta]>u]}.$
Moreover, a buyer’s average payoff across all active submarkets is given by
$\int_{\Delta(\Theta)}m\left(g\left(\frac{\mathbb{E}_{\nu}[\lambda(\theta)\theta]}{u}\right)\right)\mathbbm{1}_{[\beta\,\mathbb{E}_{\nu}[\lambda(\theta)\theta]>u]}\mathbb{E}_{\nu}[\lambda(\theta)\theta]\sigma_{\mu}(d\nu),$
@@ 5.1. Hosios condition
Despite the potential inefficiencies in the equilibrium allocation of buyers, this section identifies necessary and sufficient conditions under which the first-best outcome, as established in Proposition 1 and Proposition 2, can be decentralized. To state the result, let
$\varepsilon(t)\coloneqq\frac{m^{\prime}(t)}{m(t)}\,t$
denote the elasticity of the meeting function for $t\geq 0$.
@@ 5.2. Constrained-efficient market segmentation
This optimization problem is inherently complex: it involves optimizing over an infinite-dimensional choice variable that enters the objective function in a highly non-linear way. Specifically, the choice of segmentation influences buyers’ beliefs about seller types within each submarket, which in turn affects how buyers allocate themselves across submarkets. [^10] Although (5.2) resembles an information design problem, it differs in a crucial way: whereas a receiver’s optimal action in standard information design depends solely on the realized posterior beliefs, buyers’ search strategies here depend both on the entire distribution of posterior beliefs. In Section A.2, I derive a sufficient and necessary condition that decouples these two interdependent effects. However, even this requires solving a constrained concavification problem (Kamenica and Gentzkow, 2011) with a continuum of states. As such, a general characterization of the constrained-efficient market segmentation appears intractable in many cases.
@@
To that end, let $F$ denote the cumulative distribution function (CDF) associated with the prior $\mu\in\Delta(\Theta)$. Importantly, $F$ is also the posterior-mean distribution induced by the perfect market segmentation. [^11] Formally, $F(x)\coloneqq\sigma^{PS}_{\mu}\big{(}\{\nu\in\Delta(\Theta):\mathbb{E}_{\nu}[\theta]\leq x\}\big{)}$ for all $x\in\Theta$. More generally, a CDF $H:\Theta\to[0,1]$ represents a posterior-mean distribution induced by some market segmentation if and only if $H$ is a mean-preserving contraction of $F$ (Blackwell, 1953; Blackwell and Girshick, 1954). Let $\operatorname*{MPC}(F)$ denote the set of all such mean-preserving contractions.
@@
A segmentation $H\in\operatorname*{MPC}(F)$ is constrained-efficient if and only if
$H\in\operatorname*{arg\,max}_{\widehat{H}\in\operatorname*{MPC}(F)}\hskip 5.0ptk\int_{\Theta}\tau^{*}_{H}(x)d\widehat{H}(x).$
Furthermore,
  1. $()$
  2. $()$
@@
    is a constrained-efficient segmentation. In this case, $\tau_{H}^{*}(\underline{x})=0<\tau_{H}^{*}(\bar{x})$.
@@ A.1. Proofs
$\psi(\nu,\eta)\coloneqq\left\{\begin{array}[]{ccc}0&\mbox{if}&\beta\,\mathbb{E}_{\nu}[\theta]\leq\eta\\ f\left(\displaystyle\frac{\mathbb{E}_{\nu}[\theta]}{\eta}\right)&\mbox{if}&\beta\,\mathbb{E}_{\nu}[\theta]>\eta\end{array}\right.$.
@@ 4. $(iv)$
[…]
@@
[…]
$k\cdot\theta\int_{\Delta(\Theta)}m(\tau^{FB}_{\sigma}(\nu))\sigma(d\nu,\theta)$ $\leq k\cdot\theta\cdot m\left(\int_{\Delta(\Theta)}\tau^{FB}_{\sigma}(\nu)\sigma(d\nu,\theta)\right)$
$=k\cdot\theta\cdot m(\tilde{\tau}(\delta_{\theta}))$
$=k\cdot\theta\int_{\Delta(\Theta)}m(\tilde{\tau}(\nu))\sigma^{PS}(d\nu,\theta),$
where the inequality follows from the concavity of $m$, the first equality follows by definition of $\tilde{\tau}$, and the last follows by the definition of $\sigma^{PS}$.
[…]
@@
where the equality follows from (b), and the inequality follows from the fact $\beta$ is the highest probability with which a buyer meets a seller.
@@
where the inequality follows from (b), and the equality follows by definition that $m(0)/0=\beta$. Thus, from the contrapositive, $\beta\,\mathbb{E}_{\nu}[\lambda(\theta)\theta]>u^{*}_{\sigma}$ implies $\tau^{*}_{\sigma}(\nu)>0$. ∎
Let us now turn to proving the existence of an essentially-unique equilibrium. By Lemma 1, any search equilibrium $(\tau^{*}_{\sigma},u^{*}_{\sigma})$ satisfies $\tau^{*}_{\sigma}(\nu)=0$ for $\sigma_{\mu}$-almost all $\nu$ such that $\beta\,\mathbb{E}_{\nu}[\lambda(\theta)\theta]\leq u^{*}_{\sigma}$, and $\tau^{*}_{\sigma}(\nu)>0$ for $\sigma_{\mu}$-almost all $\nu$ such that $\beta\,\mathbb{E}_{\nu}[\lambda(\theta)\theta]>u^{*}_{\sigma}$. In the latter case, (b) implies that
@@
which from the definition of $g$ implies that $\tau^{*}_{\sigma}(\nu)=g\left(\mathbb{E}_{\nu}[\lambda(\theta)\theta]/u^{*}_{\sigma}\right)$. In other words, there exists a mapping $\phi:\Delta(\Theta)\times(0,\infty)\to\mathbb{R}_{+}$ given by
$\phi(\nu,u)\coloneqq\left\{\begin{array}[]{ccc}0&\mbox{if}&\beta\,\mathbb{E}_{\nu}[\lambda(\theta)\theta]\leq u\\[4.0pt] \displaystyle g\left(\frac{\mathbb{E}_{\nu}[\lambda(\theta)\theta]}{u}\right)&\mbox{if}&\beta\,\mathbb{E}_{\nu}[\lambda(\theta)\theta]>u\end{array}\right$.
@@
Next, observe that a pair $(\tau^{*}_{\sigma},u^{*}_{\sigma})$ jointly satisfies (b) and (4) if and only if $u^{*}_{\sigma}$ is a fixed point of the function $\Phi:(0,\infty)\to\mathbb{R}_{+}$ given by
@@
Thus far, we have shown that $\Phi(u)$ is a continuous and (weakly) decreasing function with $\lim_{u\to\beta\sup_{\theta\in\Theta}\lambda(\theta)\theta}\Phi(u)=0<\lim_{u\to 0}\Phi(u)$. Therefore, it has a unique fixed point in the interval $(0,\beta\sup_{\theta\in\Theta}\lambda(\theta)\theta)$. We have thus established the existence of an essentially-unique pair $(\tau^{*}_{\sigma},u^{*}_{\sigma})$ with $u^{*}_{\sigma}=\Phi(u^{*}_{\sigma})$ and $\tau^{*}_{\sigma}(\cdot)=\phi(\cdot,u^{*}_{\sigma})$, such that $(\tau^{*}_{\sigma},u^{*}_{\sigma})$ jointly satisfies (b) and (4). In other words, there exists an essentially-unique search equilibrium $(\tau^{*}_{\sigma},u^{*}_{\sigma})$ for each market segmentation $\sigma$. ∎
@@
Thus, for all $\theta\leq\eta_{\sigma^{PS}}/\beta$,
$\lambda(\theta)\theta\leq\lambda\left(\frac{\eta_{\sigma^{PS}}}{\beta}\right)\frac{\eta_{\sigma^{PS}}}{\beta}.$
@@
(If direction): Suppose the surplus-splitting function, $\lambda$, satisfies the properties given in the proposition. We first show that $(\tau^{FB}_{\sigma^{PS}},\lambda({\eta_{\sigma^{PS}}}/{\beta})\cdot\eta_{\sigma^{PS}})$ is a search equilibrium of $\sigma^{PS}$.
@@
where the first equality follows from the characterization of $\tau^{FB}_{\sigma^{PS}}(\delta_{\theta})$ for $\theta>\eta_{\sigma^{PS}}/\beta$ (Proposition 1), the second equality follows by the assumption on $\lambda$ given in the proposition, and the last equality follows from the fact that $f$ is the inverse of the mapping $t\mapsto 1/m^{\prime}(t)$. Therefore, the pair $(\tau^{FB}_{\sigma^{PS}},\lambda(\eta_{\sigma^{PS}})\eta_{\sigma^{PS}})$ satisfies (b) of Definition 1.